OpenStudy (cggurumanjunath):

\int { mg\cos { \theta } d\theta } =\int { mR\ddot { \theta } } d\theta

OpenStudy (cggurumanjunath):

\[\int\limits { mg\cos { \theta } d\theta } =\int\limits { mR\ddot { \theta } } d\theta \]

OpenStudy (cggurumanjunath):

integrate wrt theta ,theta dot and theta double dot. i will give medal .can any body help.

OpenStudy (vincent-lyon.fr):

Do you have a figure that goes with this equation?

OpenStudy (cggurumanjunath):

NO IT IS POLAR COORDINATE WITH CIRCULAR SURFACE,THESE ARE EQUATIONS OF MOTIONS. @Vincent-Lyon.Fr

OpenStudy (vincent-lyon.fr):

So it is a circular path. The plane is probably vertical. Is theta defined with respect the the vertical or the horizontal?

OpenStudy (cggurumanjunath):

theta is inclined angle.

OpenStudy (cggurumanjunath):

theta is inclined angle.

OpenStudy (cggurumanjunath):

@Vincent-Lyon.Fr

OpenStudy (vincent-lyon.fr):

This will work only if figure is as such: and if system is conservative. Use ΔKE = – ΔPE

OpenStudy (cggurumanjunath):

can u explain me the steps ?

OpenStudy (vincent-lyon.fr):

Express KE wrt \(\dot \theta\) Express PE wrt \(\theta\) Differentiate both and equate.

OpenStudy (cggurumanjunath):

can u explain ?

OpenStudy (vincent-lyon.fr):

Well, I cannot do it for you. Follow the steps and show your work, then we can help further.

OpenStudy (cggurumanjunath):

ke wrt theta ? what is the equation ?